# MSDformer: Multi-scale Discrete Transformer For Time Series Generation

### v1

@@ Abstract
Discrete Token Modeling (DTM), which employs vector quantization techniques, has demonstrated remarkable success in modeling non-natural language modalities, particularly in time series generation. While our prior work SDformer established the first DTM-based framework to achieve state-of-the-art performance in this domain, two critical limitations persist in existing DTM approaches: 1) their inability to capture multi-scale temporal patterns inherent to complex time series data, and 2) the absence of theoretical foundations to guide model optimization. To address these challenges, we proposes a novel multi-scale DTM-based time series generation method, called Multi-Scale Discrete Transformer (MSDformer). MSDformer employs a multi-scale time series tokenizer to learn discrete token representations at multiple scales, which jointly characterize the complex nature of time series data. Subsequently, MSDformer applies a multi-scale autoregressive token modeling technique to capture the multi-scale patterns of time series within the discrete latent space. Theoretically, we validate the effectiveness of the DTM method and the rationality of MSDformer through the rate-distortion theorem. Comprehensive experiments demonstrate that MSDformer significantly outperforms state-of-the-art methods. Both theoretical analysis and experimental results demonstrate that incorporating multi-scale information and modeling multi-scale patterns can substantially enhance the quality of generated time series in DTM-based approaches. The code will be released upon acceptance.
@@ I Introduction
To address the aforementioned challenges, we propose a novel DTM-based framework for time series generation, termed the Multi-Scale Discrete Transformer (MSDformer). This framework extends SDformer [39], advancing from single-scale to multi-scale modeling. Similar to existing DTM-based approaches, MSDformer operates in a two-stage pipeline. In the first stage, MSDformer constructs a multi-scale time series tokenizer through cascaded residual VQ-VAEs [41, 42], progressively learning discrete token representations at multiple scales. These multi-scale token sequences collectively capture the comprehensive characteristics of the original time series, enabling a more nuanced representation of both low-frequency and high-frequency features. In the second stage, MSDformer adopts an autoregressive token modeling technique to learn the distribution of multi-scale token sequences, thereby effectively capturing the multi-scale temporal patterns inherent in time series data. This technique is naturally better suited to sequential data than the masked token modeling technique, as illustrated by SDformer. Furthermore, we prove that DTM-based methods offer superior controllability in the rate-distortion trade-off compared to other continuous models, leveraging Shannon’s Rate-Distortion Theorem. Based on this, we further theoretically demonstrate that multi-scale modeling can achieve lower distortion.
In summary, our contributions include:
  • We introduce a novel DTM-based method, MSDformer, which extends SDformer by an advancement from single-scale to multi-scale modeling. This advancement enables MSDformer to effectively capture the multi-scale temporal patterns inherent in time series data, thereby enhancing generative capabilities.
  • We theoretically demonstrate that DTM-based methods offer superior controllability in the rate-distortion trade-off compared to other continuous models, and that multi-scale modeling allows for lower distortion.
  • Our experimental results confirm that the MSDformer’s performance in time series generation notably surpasses those of current methods, exemplified by an average enhancement of 34.5% in Discriminative Score and 12.6% in Context-FID Score to SDformer under long-term time series generation tasks.
II Related Work
@@
II-A Time Series Generation
@@
With the continuous advancement of deep generative models, the field of time series generation has achieved remarkable progress. The initial wave of research predominantly focuses on Generative Adversarial Networks (GANs) as the primary architecture for time series generation [15, 16, 17, 18, 19, 20]. For example, TimeGAN [17] introduces novel enhancements by incorporating an embedding function and supervised loss to better capture temporal dynamics. COT-GAN [18] innovatively combines the underlying principles of GANs with Causal Optimal Transfer theory, enabling stable generation of both low- and high-dimensional time series data. Despite these advancements, inherent challenges such as training instability and mode collapse persist, prompting researchers to explore alternative architectures.
@@
The limitations of GAN-based approaches prompt researchers to turn to Variational Autoencoders (VAEs) for time series generation. For example, TimeVAE [22], with its interpretable temporal structure, stands out as a significant advancement, delivering superior performance in time series synthesis. KoVAE [23] integrates Koopman operator theory with VAEs to provide a unified modeling framework for both regular and irregular time series data. However, these methods often face the issue of generating samples that lack detail and diversity.
@@
Denoising Diffusion Probabilistic Models (DDPMs) [43] mark a transformative phase in generative modeling, with rapid adaptation to time series generation. Early applications like DiffWave [24] demonstrate the potential of DDPMs in waveform generation, while more recent innovations such as DiffTime [25] leverages advanced score-based diffusion models for enhanced temporal data synthesis. The field reaches new heights with Diffusion-TS [26], a diffusion model framework that generates high-quality multivariate time series samples using an encoder-decoder Transformer architecture with disentangled temporal representations. While these methods exhibit superior generative capabilities, they are confronted with several challenges, including architectural complexity, limited controllability, and computational inefficiency.
II-B Discrete Token Modeling
Discrete Token Modeling (DTM) is a technique for modeling discrete token sequences using sequence models, originally widely applied in natural language modalities. In recent years, the rise of large-scale language models has driven a new era of development in artificial intelligence [27, 28], highlighting the efficiency of the DTM technique. Building on its success in natural language tasks, DTM has been extended to non-natural language modalities—such as vision, audio, and time series—through vector quantization models like VQ-VAE [42] and VQ-GAN [44].
Early DTM approaches in non-natural language modalities primarily rely on two paradigms: Autoregressive Token Modeling (ARTM) and Masked Token Modeling (MTM). ARTM predicts the next token in a sequence based on previous tokens using a categorical distribution. Notable examples include DALL-E [29] and Parti [30] for text-to-image generation, and T2M-GPT [45] and MotionGPT [46] for text-to-motion synthesis. In contrast, MTM employs a masked token prediction objective [47], reconstructing randomly masked tokens using the observed context. This approach has been adopted by models like MaskGIT [31] and MUSE [32] for image generation, and extended to video generation in MAGVIT [33] and Phenaki [34], demonstrating its versatility.
Recently, a novel DTM paradigm called Visual AutoRegressive Modeling (VAR)[48] has emerged. VAR redefines autoregressive modeling as coarse-to-fine ”next-scale prediction” or ”next-resolution prediction”. During tokenization, images are encoded into multi-scale token maps via hierarchical quantization. The generation process then proceeds autoregressively from low to high resolutions, with each step predicting the next higher-resolution token map. This multi-scale autoregressive approach enhances spatial coherence and preserves local details, resulting in higher-quality and more natural image synthesis.
@@ II-B Discrete Token Modeling
III Problem Statement
@@
IV Methods
Figure 1: The workflow of multi-scale time series tokenizer. The multi-scale time series tokenizer comprises $K$ modules, each of which is a VQ-VAE model that employs a similarity-driven vector quantization method to learn discrete token representations at different scales. Specifically, the first scale is used to model the large-scale information of the time series, while each subsequent scale models the difference between the input and output of the previous scale. The discrete token sequences obtained at each scale are then combined to jointly characterize the complete time series data. Notably, when $K=1$, the model degenerates to SDformer’s single-scale time series tokenizer [39], as shown by the red dashed line in the figure.
Figure 2: The workflow of the second stage of MSDformer: Multi-Scale Autoregressive Token Modeling. In this stage, the [BOS] token is first concatenated with the discrete token sequence of each scale. The resulting sequence, along with the token type ID and position ID, is then fed into the Decoder-only Transformer to predict the next token at each position. Notably, when $K=1$, the model degenerates to SDformer’s single-scale autoregressive token modeling [39].
@@ IV Methods
IV-A Multi-scale Time Series Tokenizer
The multi-scale tokenizer implements cascaded residual learning [41] through stacked VQ-VAE models [42], where each subsequent model hierarchically encodes residual information from the preceding scale. This coarse-to-fine decomposition progressively refines temporal features, yielding more precise and detailed representations of the underlying time series data.
@@ IV-A Multi-scale Time Series Tokenizer
(2) Vector Quantization. Then, the similarity-driven vector quantization $Q(\ \cdot\ ;C^{(k)})$ identifies the index of the vector in the codebook $C^{(k)}$ that exhibits the highest similarity to $h_{i}^{(k)}$, which can be expressed as:
@@
where the first loss is the reconstruction loss, the second loss is the embedding loss, sg()($\cdot$)( ⋅ ) represents the stop gradient, and $\lambda$ is hyperparameter used to adjust the weight of different parts.
@@
IV-B Multi-scale Autoregressive Token Modeling
In this subsection, we focus on utilizing an autoregressive token modeling technique to learn the data distribution of the time series in the multi-scale discrete token space.
@@ IV-B Multi-scale Autoregressive Token Modeling
where $\mathbf{E}_{s}\in\mathbb{R}^{(K+1)\times d_{m}}$ is the token type encoding matrix, $K$ is the number of scales, and the additional dimension is reserved for the [BOS] token. Here, $k_{i}\in\{0,1,\ldots,K\}$ is the scale identifier for the token, defined as follows.
@@
Specifically, the [BOS] token corresponds to $k_{i}=0$, while the $k_{i}$ for other tokens is uniquely determined by the scale interval $[T^{(k-1)}+1,T^{(k)}]$ in which their index $i$ falls. The corresponding token type encoding sequence structure can be represented as:
@@
IV-C Inference Stage of MSDformer
@@
9:$V\leftarrow\sum_{k=1}^{K}V^{(k)}$triangleright Total vocabulary size
10:$L\leftarrow\sum_{k=1}^{K}\tau/r^{(k)}$triangleright Total sequence length
11:$\hat{y}_{0}\leftarrow V$triangleright Initialize with BOS token
@@
13:     $\mathcal{Z}\leftarrow\mathcal{G}_{ms}(\hat{y}_{0:t-1})$triangleright Generate token logits
14:     $p\leftarrow\text{Softmax}(\mathcal{Z}[-1])$triangleright Compute probabilities
15:     $\mathcal{I}\leftarrow\text{Sample}(p)$triangleright Sample next token
16:     $\hat{y}_{0:t}\leftarrow\hat{y}_{0:t-1}\oplus\mathcal{I}$triangleright Update sequence
@@
19:$i\leftarrow 1$triangleright Initialize pointe
20:$S^{(0)}\leftarrow 0$triangleright cumulative vocabulary size
@@
22:     $S^{(k)}\leftarrow S^{(k-1)}+V^{(k)}$triangleright Update vocabulary offset
23:     $L^{(k)}\leftarrow\tau/r^{(k)}$triangleright Calculate scale-specific length
24:     $\hat{y}^{(k)}_{1:L^{(k)}}\leftarrow\hat{y}_{i:i+L^{(k)}}-S^{(k-1)}$triangleright Allocate tokens
25:     $i\leftarrow i+L^{(k)}$triangleright Move pointer
@@
30:     $\hat{f}_{1:\tau}^{(k)}\leftarrow\mathcal{D}^{(k)}(\hat{h}_{1:L^{(k)}}^{(k)})$triangleright Decoding
31:end for
32:$\hat{x}_{1:\tau}\leftarrow\sum_{k=1}^{K}\hat{f}_{1:\tau}^{(k)}$triangleright Aggregate multi-scale outputs
@@
V Theoretical Analysis
V-A Theoretical Analysis of Discrete Token Modeling

### v2

@@ Abstract
Discrete Token Modeling (DTM), which employs vector quantization techniques, has demonstrated remarkable success in modeling non-natural language modalities, particularly in time series generation. While our prior work SDformer established the first DTM-based framework to achieve state-of-the-art performance in this domain, two critical limitations persist in existing DTM approaches: 1) their inability to capture multi-scale temporal patterns inherent to complex time series data, and 2) the absence of theoretical foundations to guide model optimization. To address these challenges, we proposes a novel multi-scale DTM-based time series generation method, called Multi-Scale Discrete Transformer (MSDformer). MSDformer employs a multi-scale time series tokenizer to learn discrete token representations at multiple scales, which jointly characterize the complex nature of time series data. Subsequently, MSDformer applies a multi-scale autoregressive token modeling technique to capture the multi-scale patterns of time series within the discrete latent space. Theoretically, we validate the effectiveness of the DTM method and the rationality of MSDformer through the rate-distortion theorem. Comprehensive experiments demonstrate that MSDformer significantly outperforms state-of-the-art methods. Both theoretical analysis and experimental results demonstrate that incorporating multi-scale information and modeling multi-scale patterns can substantially enhance the quality of generated time series in DTM-based approaches.
@@ 1 Introduction
To address the aforementioned challenges, we propose a novel DTM-based framework for time series generation, termed the Multi-Scale Discrete Transformer (MSDformer). This framework extends SDformer [67], advancing from single-scale to multi-scale modeling. Similar to existing DTM-based approaches, MSDformer operates in a two-stage pipeline. In the first stage, MSDformer constructs a multi-scale time series tokenizer through cascaded residual VQ-VAEs [36, 48], progressively learning discrete token representations at multiple scales. These multi-scale token sequences collectively capture the comprehensive characteristics of the original time series, enabling a more nuanced representation of both low-frequency and high-frequency features. In the second stage, MSDformer adopts an autoregressive token modeling technique to learn the distribution of multi-scale token sequences, thereby effectively capturing the multi-scale temporal patterns inherent in time series data. This technique is naturally better suited to sequential data than the masked token modeling technique, as illustrated by SDformer. Furthermore, we prove that DTM-based methods offer superior controllability in the rate-distortion trade-off compared to other continuous models, leveraging Shannon’s Rate-Distortion Theorem. Based on this, we further theoretically demonstrate that multi-scale modeling can achieve lower distortion. In summary, our contributions include:
  • Multi-scale Generation. We introduce a novel DTM-based method, MSDformer, which extends SDformer by an advancement from single-scale to multi-scale modeling. This advancement enables MSDformer to effectively capture the multi-scale temporal patterns inherent in time series data, thereby enhancing generative capabilities.
  • Strong Controllability. We develop a rate–distortion analysis theoretically demonstrating that discrete tokenization induces an explicit, adjustable rate constraint via finite codebooks, thereby providing stronger controllability than continuous latent-variable models. We further show that, under the same overall rate budget, allocating capacity across scales can reduce distortion more effectively than increasing the codebook size in a single-scale model.
  • Long-term Generation Capability. Extensive experiments demonstrate that MSDformer consistently outperforms SDformer and other continuous generative models, with particularly clear gains in long-term time-series generation. Compared with SDformer, MSDformer achieves average improvements of 34.5% in Discriminative Score and 12.6% in Context-FID.
2 Related Work
@@
2.1 Time Series Generation
@@
Generative Adversarial Networks (GANs)-based Models. With the continuous advancement of deep generative models, the field of time series generation has achieved remarkable progress. The initial wave of research predominantly focuses on Generative Adversarial Networks as the primary architecture for time series generation [34, 19, 59, 56, 37, 23]. For example, TimeGAN [59] introduces novel enhancements by incorporating an embedding function and supervised loss to better capture temporal dynamics. COT-GAN [56] innovatively combines the underlying principles of GANs with Causal Optimal Transfer theory, enabling stable generation of both low- and high-dimensional time series data.
GANs-based models, despite their progress, remain notoriously hard to train due to optimization instability [33, 26, 44]. They often suffer from non-convergence (oscillatory training), vanishing gradients when the discriminator becomes overly strong, and mode collapse [45]. The latter is especially detrimental for time-series generation [4, 52], where the model may degenerate to producing a small set of highly similar sequences (e.g., repetitive periodic patterns), failing to capture the diversity and fine-grained temporal dynamics of real data.
@@
Variational Autoencoders (VAEs)-based Models. Since VAEs are are known to be more robust to the GANs-based issues above, some attentions are turned to this type of generative paradigm for time series generation. For example, TimeVAE [13], with its interpretable temporal structure, stands out as a significant advancement, delivering superior performance in time series synthesis. KoVAE [35] integrates Koopman operator theory with VAEs to provide a unified modeling framework for both regular and irregular time series data.
However, traditional VAEs for time-series generation typically rely on either an unstructured Gaussian prior or structured (e.g., approximately linear) latent dynamics. While these designs can aid inference or dynamical interpretability, they do not necessarily impose explicit constraints aimed at latent controllability, so the learned representations may remain entangled and hard to manipulate. Consequently, controllable generation is limited, and posterior collapse can still occur.
@@
Denoising Diffusion Probabilistic Models (DDPMs). Denoising Diffusion Probabilistic Models [22] mark a transformative phase in generative modeling, with rapid adaptation to time series generation. Early applications like DiffWave [28] demonstrate the potential of DDPMs in waveform generation, while more recent innovations such as DiffTime [11] leverages advanced score-based diffusion models for enhanced temporal data synthesis. The field reaches new heights with Diffusion-TS [62], a diffusion model framework that generates high-quality multivariate time series samples using an encoder-decoder Transformer architecture with disentangled temporal representations.
While these approaches achieve competitive results, diffusion models typically map non-i.i.d. time-series distributions to an isotropic Gaussian prior via iterative noise injection, producing long and complex trajectories. This may hinder sampling efficiency and limit performance, especially in long-term generation setting, which are shown in our following experimental section.
2.2 Discrete Token Modeling
In contrast to continuous generative models above, which operate in a non-explicit constrained latent space or require costly iterative sampling, discrete token modeling strategy provides a new generation paradigm. Discrete Token Modeling (DTM) is a technique for modeling discrete token sequences using sequence models, originally widely applied in natural language modalities [15]. In comparison, GAN- and VAE-based generators often struggle to provide reliable and fine-grained control: GANs are prone to unstable training and mode collapse, while VAEs may yield entangled latents and suffer from posterior collapse, limiting controllable synthesis. Diffusion models improve sample quality but typically require dozens to hundreds of denoising steps, which makes long-horizon generation computationally expensive. By discretizing time series into an explicit token space, DTM offers a more controllable and efficient alternative, where generation is performed in an explicit, finite code space that is easier to constrain, manipulate, and sample efficiently. In recent years, the rise of large-scale language models has driven a new era of development in artificial intelligence [1, 3], highlighting the efficiency and effectiveness of the DTM technique in sequence data. Moreover, DTM has been extended to non-natural language modalities—such as iamge [40], video [57], and audio [17] through vector quantization models like VQ-VAE [48] and VQ-GAN [18].
Early discrete token modeling (DTM) in non-language modalities mainly follows two paradigms: Autoregressive Token Modeling (ARTM) and Masked Token Modeling (MTM). ARTM predicts the next token from previous tokens using a categorical distribution, and has been widely used in text-to-image generation (e.g., DALL-E [40], Parti [60]) as well as text-to-motion synthesis (e.g., T2M-GPT [65], MotionGPT [25]). In contrast, MTM adopts a masked prediction objective [14], reconstructing randomly masked tokens from the observed context, which underlies models such as MaskGIT [7] and MUSE [6] for efficient and effective image generation. More recently, Visual AutoRegressive Modeling (VAR) [46] introduces a new DTM paradigm that reframes autoregression as coarse-to-fine “next-scale ” prediction: images are tokenized into multi-scale token maps via hierarchical quantization, and generation proceeds autoregressively from low to high resolutions by predicting the next higher-resolution token map, improving global coherence while preserving fine local details. Also, this type of multi-scale modeling perspective is also naturally compatible with time-series data [8, 9, 20], where temporal dynamics always exhibit hierarchical structure.
@@ 2.2 Discrete Token Modeling
However, our previous SDformer adopts a single-scale, similarity-based discrete modeling strategy in the latent space. While effective, it overlooks the inherent multi-scale nature of time-series data and lacks a systematic theoretical and empirical investigation of why discrete modeling is beneficial. In this work, we demonstrate that MSDformer consistently improves short- and long-term generation, outperforming SDformer as well as representative continuous generative baselines. Moreover, we find that multi-scale token modeling yields larger gains in generation quality than simply increasing the codebook size at a single scale, highlighting the advantage of explicitly modeling temporal hierarchies.
3 Problem Statement
@@
4 Methods
Figure 1: Workflow of the multi-scale time-series tokenizer. The tokenizer consists of (K) modules, each implemented as a VQ-VAE with similarity-driven vector quantization to learn discrete tokens at a specific scale. The first module captures the coarse, large-scale structure of the series, while each subsequent module models the residual (i.e., the difference between the input and the reconstruction from the previous scale). Token sequences from all scales are then aggregated to represent the full time series. When (K=1), the tokenizer reduces to the SDformer [67], indicated by the red dashed line.
Figure 2: Stage-2 workflow of MSDformer (multi-scale autoregressive token modeling). A [BOS] token is prepended to the concatenation result of each scale’s discrete token sequence, and the resulting tokens with type and position IDs are fed into a decoder-only Transformer to autoregressively predict the next token. When (K=1), it reduces to ourSDformer’s single-scale token modeling [67].
@@ 4 Methods
4.1 Multi-scale Time Series Tokenizer
The multi-scale tokenizer implements cascaded residual learning [36] through stacked VQ-VAE models [48], where each subsequent model hierarchically encodes residual information from the preceding scale. This coarse-to-fine decomposition progressively refines temporal features, yielding more precise and detailed representations of the underlying time series data. Algorithm 1 outlines the training procedure for the multi-scale time series tokenizer in this stage, in which our objective is to train a multi-scale time series tokenizer that can produce high-quality discrete representations of time series data while ensuring excellent reconstruction performance.
@@ 4.1 Multi-scale Time Series Tokenizer
[…]
@@
where the first loss is the reconstruction loss, the second loss is the embedding loss, sg()($\cdot$) represents the stop gradient, and $\lambda$ is hyperparameter used to adjust the weight of different parts.
@@
Algorithm 1 Training of multi-scale time series tokenizer.
Input: Time series dataset $D=\{x_{1:\tau}^{i}\}_{i=1}^{n}$
1:for $i\leftarrow 1$ to $N_{iter}$ do
2:  Get the $X_{1:\tau}$ $\sim$ $D$;
3:  Initialize residual input $f_{1:\tau}^{(1)}\leftarrow X_{1:\tau}$;
4:  for $k\leftarrow 1$ to $K$ do
5:   Get the $h_{1:L^{(k)}}^{(k)}$ by Equation (2);
6:   Get the $y_{1:L^{(k)}}^{(k)}$ by Equation (3);
7:   Get the $\tilde{h}_{1:L^{(k)}}^{(k)}$ by Equation (4);
8:   Get the $\tilde{f}_{1:\tau}^{(k)}$ by Equation (5);
9:   Get the $f_{1:\tau}^{(k+1)}$ by Equation (6);
10:  end for
11:  Get the $\tilde{x}_{1:\tau}$ by Equation (7);
12:  Compute the training loss $\mathcal{L}$ by Equation (8);
14:  Update the codebook by Equations (9) and (11);
15:end for
4.2 Multi-scale Autoregressive Token Modeling
In this subsection, we focus on utilizing an autoregressive token modeling technique to learn the data distribution of the time series in the multi-scale discrete token space. In this stage, we aim to train a multi-scale autoregressive discrete Transformer that learns the distribution of multi-scale discrete token sequences, and the training process is shown in Algorithm 2. Besides, for the training algorithms for our SDformer, simply set the number of scales $K$ in MSDformer to 1.
@@ 4.2 Multi-scale Autoregressive Token Modeling
1:for $i\leftarrow 1$ to $N_{iter}$ do
3:  Get the label $y_{1:L}$ by Equations (12) and (13);
4:  Get the input $\tilde{y}_{1:L}^{in}$ by Equations (14) and (15);
5:  Feed the $\tilde{y}_{1:L}^{in}$ to $\mathcal{G}_{ms}$;
6:  Compute the training loss $\mathcal{L}_{ms}$ by Equation (22);
8:end for
9:Return trained $\mathcal{G}_{ms}$.
@@
where $\mathbf{E}_{s}\in\mathbb{R}^{(K+1)\times d_{m}}$ is the token type encoding matrix, which uses to explicitly indicate which scale each token comes from. Since tokens from different scales are concatenated into a single input sequence, this embedding prevents the Transformer from confusing tokens of different temporal resolutions and enables scale-aware attention and prediction, $K$ is the number of scales, and the additional dimension is reserved for the [BOS] token. Here, $k_{i}\in\{0,1,\ldots,K\}$ is the scale identifier for the token, defined as follows.
@@
Specifically, the [BOS] token corresponds to $k_{i}=0$, while the $k_{i}$ for other tokens is uniquely determined by the scale interval $T^{(k-1)}+1,T^{(k)}]$ in which their index $i$ falls. The corresponding token type encoding sequence structure can be represented as:
@@
4.3 Inference Stage of MSDformer
@@
[…]
12:for $t=1$ toLL do
13:  $\mathcal{Z}\leftarrow\mathcal{G}_{ms}(\hat{y}_{0:t-1})$ $\triangleright$ Generate token logits
14:  $p\leftarrow\text{Softmax}(\mathcal{Z}[-1])$ $\triangleright$ Compute probabilities
15:  $\mathcal{I}\leftarrow\text{Sample}(p)$ $\triangleright$ Sample next token
16:  $\hat{y}_{0:t}\leftarrow\hat{y}_{0:t-1}\oplus\mathcal{I}$ $\triangleright$ Update sequence
@@
[…]
21:for $k=1$ toKK do
22:  $S^{(k)}\leftarrow S^{(k-1)}+V^{(k)}$ $\triangleright$ Update vocabulary offset
23:  $L^{(k)}\leftarrow\tau/r^{(k)}$ $\triangleright$ Calculate scale-specific length
24:  $\hat{y}^{(k)}_{1:L^{(k)}}\leftarrow\hat{y}_{i:i+L^{(k)}}-S^{(k-1)}$ $\triangleright$ Allocate tokens
25:  $i\leftarrow i+L^{(k)}$ $\triangleright$ Move pointer
@@
5 Theoretical Analysis
5.1 Theoretical Analysis of Discrete Token Modeling